OpenStudy (anonymous):

Please help me solve each equation: a) 3x-2/x+4 = 0 b) 2/x-3 = 0

hero (hero):

@rcatabijan, you here?

OpenStudy (anonymous):

@hero, im here

OpenStudy (kc_kennylau):

Are there any parentheses missing?

OpenStudy (anonymous):

@kc_kennylau nope, thats it thats shown

OpenStudy (kc_kennylau):

Sometimes parentheses are necessary when converting things from Mathematical notation to text. For example, you can't write \(\dfrac{x-2}3\) as x-2/3, but you have to add parentheses so that it becomes (x-2)/3 for ambiguity.

OpenStudy (kc_kennylau):

Let me ask you again, are there any parentheses missing?

OpenStudy (anonymous):

Yes, for a) I believe that you're missing parenthesis for the top and bottom in order to convert into mathematical notations yes? (3x-2)/(x+4)=0?

OpenStudy (kc_kennylau):

Do you have any idea?

OpenStudy (kc_kennylau):

For example, if you know that A/B=0, what can we say about the value of A?

OpenStudy (anonymous):

would we need to isolate "A" in order to get the answer?

OpenStudy (kc_kennylau):

Yep, that would be necessary.

OpenStudy (anonymous):

Can I move (3x-2) to the other side? would that be the first step? >< or can I divide the like terms?

OpenStudy (kc_kennylau):

The first step would be to multiply both sides by the denominator.

OpenStudy (anonymous):

3x^2-8 / x^2 + 16

OpenStudy (anonymous):

right?

OpenStudy (kc_kennylau):

No, that is not correct. Firstly, (3x-2)(x+4) is not 3x^2-8, you cannot multiply the first terms together and the last terms together and then add them up. Secondly, (x+4)(x+4) is not x^2+16, with the exact same reason as above. Thirdly, I was referring to the both sides of the equation when I was saying "sides".

OpenStudy (anonymous):

Can you possibly guide me with the first question? I really don't know where to start

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} \frac{3x-2}{x+4}&=&0 \end{array}\]

OpenStudy (kc_kennylau):

Multiply both sides by x+4: \[\begin{array}{rcl} \frac{3x-2}{\color{blue}{x+4}}\times(\color{blue}{x+4})&=&0\times(\color{blue}{x+4}) \end{array}\]

OpenStudy (kc_kennylau):

What can you do now?

OpenStudy (anonymous):

cancel x+4 on the left side, then try to isolate for "x"?

OpenStudy (kc_kennylau):

What do you get after cancelling?

OpenStudy (anonymous):

3x -2 = x+4

OpenStudy (kc_kennylau):

How did you get the x+4 on the right hand side?

OpenStudy (anonymous):

because of " 0×(x+4)"

OpenStudy (anonymous):

or do you not do that?

OpenStudy (anonymous):

cause usually when you multiply one side you do to the other as well, well thats what I learned

OpenStudy (anonymous):

Im reading in my note that the numerator must equal zero but not the denominator. Therefore, 3x−2 x+4 ×(x+4) = 0×(x+4) **WIll give me: 3x-2 = 0 3x = 2 x = 2/3???

OpenStudy (kc_kennylau):

Please note that 0(x+4) is 0 instead of x+4. I notice that you have corrected this mistake. Yes, your answer is perfectly correct.

OpenStudy (anonymous):

omg thank you. One more question to go. What if I were to do the same for the second question BUT it contains the "x". would it be just 2 = 0? I assume this is incorrect.

OpenStudy (anonymous):

the denominator contains the "x"

OpenStudy (kc_kennylau):

What you have done is perfectly correct.

OpenStudy (kc_kennylau):

Think about it this way. Think of (x-3) as one number. Now you have 2 divided a number equals to 0. 2 divided by what is equal to 0?

OpenStudy (anonymous):

0?

OpenStudy (kc_kennylau):

2/0=0 ?!

OpenStudy (anonymous):

That's total wrong.. would it be undefined? because my calc says error

OpenStudy (kc_kennylau):

Correct

OpenStudy (anonymous):

it seemed like a tricky question..

OpenStudy (anonymous):

I feel better now, thanks Kc, I'm going to continue on my other work and write down my stepped answers. You're a life saver

OpenStudy (anonymous):

I'm doing MHF 4U, I may have more questions, is there any way I can contact you for more help if I need it in the future?

OpenStudy (kc_kennylau):

Here :)

OpenStudy (anonymous):

Bookmarked your page so I can message you :) Thanks again! :D